OpenStudy (anonymous):

can any one help please. consider the triangle with vertices A(1,2,3), B(1,-1,9), C(13,4,5) find AB ,AC, ABXAC, what is the area of the triangle?

OpenStudy (anonymous):

AB=B-A={0, -3, 6} AC=C-A={12, 2, 2}

OpenStudy (anonymous):

\[ AB \times AC=\{-18, 72, 36\} \] Compute the above using determinant.

OpenStudy (anonymous):

Area of the triangle ABC is half of the magnitude of AB X AC.

OpenStudy (anonymous):

\[ Area = 9 \sqrt{21} \]

OpenStudy (anonymous):

thank you but in ther AC isnt (2,2,2) how did you get it (12,2,2)

OpenStudy (anonymous):

AC =C-A=(12,2,2)